A 21-inch rope is to be cut into two pieces so that one piece is 5 inches longer than twice the other. How long is each piece, if a = 2, b = 5, and l = 21?
Answer by Maths68(1474)

Let
One Piece = x inches
Other Piece = 21-x inches
one piece is 5 inches longer than twice the other
x=2(21-x)+5
x=42-2x+5
x=47-2x
x+2x=47
3x=47
3x/3=47/3
x=15.66
One Piece = x = 15.66 inches
Other Piece = 21-x = 21-15.66 = 5.33 inches


Check
x=2(21-x)+5
15.66=2(21-15.66)+5
15.66=(42-31.32)+5
15.66=10.68+5
15.66=15.68
round it off to 1 dp
15.7=15.7